# Irreversibility and symmetry breaking in the creation and annihilation of defects in active living matter

### v1

@@ III Methods
The parameter $k$, setting the number of neighbors, should be taken to be proportional to $\sqrt{N}$ Beirlant et al. (1997). Estimates reported for single-defects, single cells and the nematic/flow fields used $k=\sqrt{N}$. EP of defect pairs use only the orientation of defect, which are discretized. As a result, angles are not unique and the distence between neighbors may be zero. To this end, we use a larger value of k=2sqrtNk=2sqrt{N}.
Combining polar and nematic vector fields
The cells are long, thin objects with a shape that is broadly identical under $\pi$ rotation of the director. At high density, they form a thin layer with local orientational order. To describe the average orientation local of the we cells we introduce the director field $\hat{n}$ defined as:
$\underline{\hat{n}}=[\cos(\psi),\sin(\psi)]$ (7)
where $\psi$ is the average orientation of the cells at a point in space.
Most observations on the cells are compatible with the typical nematic active stress:
$\sigma_{ij}\sim(\hat{n}_{i}\hat{n}_{j}-\delta_{ij}/2)$ (8)
This is the lowest order active term consistent with conservation of momentum and nematic symmetry.
However, the cells here are able to self-propel, exchanging momentum with the substrate leading to an additional active body force on the material given. We introduce the second director field $\hat{p}$, which describes the self-propulsion direction of the cells. This can lead to a body force of the form:
$\underline{f}\sim\underline{\hat{p}}$ (9)
The cells are elongated in shape and self-propel along their long axis, meaning that $\underline{\hat{p}}$ and $\underline{\hat{n}}$ must be either parallel or anti-parallel, i.e. $(\underline{\hat{p}}.\underline{\hat{n}})^{2}=1$.
We introduce a polar director angle $\theta$ such that:
$\underline{\hat{p}}=[\cos(\theta),\sin(\theta)]$ (10)
Which implies that either $\theta=\psi$ or $\theta=\psi+\pi$.
@@
Topological defects within the nematic field are characterized by a half integer winding number $k=\pm^{1}/_{2}$. The director field at a polar angle $\phi$ around a topological defect is given by:
@@
A half integer winding number is incompatible with the polar field $\hat{p}$, and necessitates the introduction of a discontinuous jump in the polar director angle, $\theta$. This is the grain boundary over which $\underline{\hat{p}}$ goes through a $\pi$ rotation which begins and ends at a topological defect. We introduce the grain boundary as follows
@@
$\underline{F}$ $=\oint\underline{f}(\phi)\textrm{d}\phi$ (14)
$\sim\oint\underline{\hat{p}}(\phi)\textrm{d}\phi$ (15)
@@
$\underline{F}$ $\sim\int_{-\pi}^{\phi_{g}}\underline{\hat{p}}(\phi)\textrm{d}\phi+\int_{\phi_{g}}^{\pi}\underline{\hat{p}}(\phi)\textrm{d}\phi$
$=\int_{-\pi}^{\phi_{g}}[\cos(k\phi+m\pi),\sin(k\phi+m\pi)]\textrm{d}\phi+$
$+\int_{\phi_{g}}^{\pi}[\cos(k\phi+(m+1)\pi),\sin(k\phi+(m+1)\pi)]\textrm{d}\phi$
$=\left[\frac{[\sin(k\phi+m\pi),-\cos(k\phi+m\pi)]}{k}\right]_{-\pi}^{\phi_{g}}+$
$+\left[\frac{[\sin(k\phi+(m+1)\pi),-\cos(k\phi+(m+1)\pi)]}{k}\right]_{\phi_{g}}^{\pi}$
$=\frac{2}{k}\left[\sin\left(k\phi_{g}+m\pi\right),-\cos\left(k\phi_{g}+m\pi\right)\right]$
$=\frac{2}{k}\left[\cos\left(k\phi_{g}+(m-0.5)\pi\right),\sin\left(k\phi_{g}+(m-0.5)\pi\right)\right]$
@@
First, there is a net translational force on both $\pm^{1}/_{2}$ defects, meaning that we expect both defects to self-propel. This is in addition to the self-propulsion force already present for $^{1}/_{2}$ defects due to the nematic active stress.
If $\phi_{g}=0$, as is the case immediately after creation, the additional force is perfectly perpendicular to the symmetry axis of the defect. This would imply a self-propulsion perpendicular to the tail for $^{1}/_{2}$ defects (or a leg for $^{1}/_{2}$ defects).
@@
We can assess the active torsional force around a defect caused by the polar active force by azimuthally integrating the azimuthal component of the polar force. This is given by:
@@
$\tau$ $\sim\int_{-\pi}^{\pi}\sin(\theta-\phi)\textrm{d}\phi$ (19)
$=\int_{-\pi}^{\pi}\sin((k-1)\phi+\pi*(H(\phi-\phi_{g})+m))\textrm{d}\phi$ (20)
$=\left[\frac{-\cos((k-1)\phi+m\pi)}{k-1}\right]_{-\pi}^{\phi_{g}}+$ (21)
$+\left[\frac{-\cos((k-1)\phi+(m+1)\pi)}{k-1}\right]_{\phi_{g}}^{\pi}$ (22)
$=\frac{1}{k-1}\left[2\cos((k-1)\phi_{g}+(m+1)\pi)+\right$. (23)
@@
Second, the magnitude of the torque depends on the orientation of the grain boundary. The torque disappears when $(k-1)\phi_{g}+m\pi=(n+0.5)\pi$, $n\in\mathbb{Z}$. This is equivalent to when the net additional net force aligns with an axis of miror symmetry.
@@
IV Supplementary material
Table 3: Entropy production (EP) estimation - control. Top: EP for trajectories of individual bacteria and cells. Bottom: EP for the nematic and flow fields at random positions. Values show the estimate $\pm$ the standard error. Since this is a lower bound, a value that is statistically larger than zero implies non reversible dynamics. See the methods section for details.
Bacteria Cells
EP individual trajectories $004\pm 0.004 0.0$$003\pm 0.0002$
EP nematic+flow $008\pm 0.001 0.000$$008\pm 0.001$

### v2

@@ III Methods
The parameter $k$, setting the number of neighbors, should be taken to be proportional to $\sqrt{N}$ Beirlant et al. (1997). Estimates reported used $k=\sqrt{N}$.
Simulating an active nemato-polar system
We start from the standard description of an incompressible two-dimensional active nematic. The nematic orientation of the cells is described by the $\bm{Q}$ tensor which combines the cellular orientation $\bm{\hat{n}}$ and the local degree of nematic order $S\in[0,1]$ into a single tensorial order parameter $Q_{ij}=S(n_{i}n_{j}-\delta_{ij}/2)$. The motion of the cells are described by the velocity field $\bm{v}$ which abides by $\nabla.\bm{v}=0$.
The dynamics of the nematic tensor are given by
$D_{t}\bm{Q}=\lambda_{Q}S\bm{u}+\frac{1}{\gamma_{Q}}\bm{H}_{Q}+\bm{Q}.\bm{\omega}-\bm{\omega}.\bm{Q},$ (8)
where $D_{t}=\partial_{t}+\bm{v}.\nabla$ is the material derivative. $\bm{u}$ and $\bm{\omega}$ represent the strain rate and vorticity tensors, respectively. The molecular tensor is given by the functional derivative of the elastic free energy with respect to the nematic tensor, $\bm{H}=-\delta F/\delta\bm{Q}$ and $\gamma_{Q}$ is the rotational viscosity associated with the nematic field. The variable $\lambda_{Q}$ describes the tendency of the nematic field to align with the flow.
To describe the polar orientation of the cells we introduce the vectorial order parameter $\bm{p}$ which describes the average local orientation of the self-propulsion direction of the cells. This follows similar dynamics to the nematic tensor, given by
$D_{t}\bm{p}=\lambda_{p}\bm{u}.\bm{p}+\frac{1}{\gamma_{p}}\bm{h}_{p}-\bm{\omega}.\bm{p}.$ (9)
Here, the molecular vector is given by the functional derivative of the elastic free energy with respect to the polar field, $\bm{h}=-\delta F/\delta\bm{p}$ and $\gamma_{p}$ is the rotational viscosity associated with the polar field. As before, the variable $\lambda_{p}$ describes the tendency of the polar field to align with the flow.
The elastic free energy governs the relaxation dynamics of the $\bm{Q}$ and $\bm{p}$ fields. We split it into 3 parts,
$F=F_{Q}+F_{C}+F_{p}.$ (10)
The first part contains the nematic energy and describes the steric interactions between cells. It is given by
$F_{Q}=\frac{1}{2}\int\Big[K_{Q}|\nabla\bm{Q}|^{2}+\alpha_{Q}\textrm{tr}\bm{Q}^{2}(\textrm{tr}\bm{Q}^{2}-1)\Big]\textrm{d}A.$ (11)
In the expression above, $K_{Q}$ denotes the single elastic constant and $\alpha_{Q}$ controls the strength of the Landau terms which apply a soft constraint pushing the system into the nematic phase $S=1$.
The second part of the energy is the coupling term between the $\bm{Q}$ and $\bm{p}$ fields given by
$F_{C}=-\frac{\alpha_{C}}{2}\bm{p}.\bm{Q}.\bm{p}.$ (12)
This energy is minimized when the $\bm{Q}$ and $\bm{p}$ fields align and is used to ensure they remain compatible descriptions of the cellular orientations.
The final part of the energy describes the tendency to locally align polarity and is given by
$F_{p}=\frac{1}{2}\int\Big[K_{p}|\nabla\bm{p}|^{2}+\frac{\alpha_{p}}{4}\bm{p}^{4}-\frac{\beta_{p}}{2}\bm{p}^{2}\textrm{d}A.$ (13)
Here, $K_{p}$ is the elastic constant associated with gradients in the polar field and $\alpha_{p}$ and $\beta_{p}$ control the transition to polar order. This energy is employed predominantly to maintain the polar order parameter and keep the field continuous, thus it has a smaller magnitude than the other energies. We typically set $\beta_{p}=0$. This implies that the polar field becomes disordered when there is no nematic order, for example at the core of a defect.
The dynamics of the velocity evolve according to
$\rho D_{t}\bm{v}=\eta\nabla^{2}\bm{v}-P\bm{I}+\nabla.\bm{\sigma}^{p}+\zeta_{Q}\nabla.\bm{Q}+\zeta_{p}\bm{p}-\mu\bm{v}.$ (14)
Here, $\rho$ is a constant density, $\eta$ is the shear viscosity, and $\mu$ is the friction coefficient. $\zeta_{Q}$ is the magnitude of the nematic active stress and $\zeta_{p}$ is the magnitude of the polar active force. $P$ if the pressure which is fixed by the incompressibility condition and $\bm{I}$ is the identity matrix.
Finally, the passive stress is given by
$\bm{\sigma}^{p}=-\lambda_{Q}S\bm{H}+\bm{Q}.\bm{H}-\bm{H}.\bm{Q}-\lambda_{p}\bm{h}\odot\bm{p}-\bm{h}\wedge\bm{p}.$ (15)
The symmetric and antisymmetric tensor products are defined as
$\bm{h}\odot\bm{p}$ $=\frac{1}{2}(h_{i}p_{j}+h_{j}p_{i})$ (16)
$\bm{h}\wedge\bm{p}$ $=\frac{1}{2}(h_{i}p_{j}-h_{j}p_{i}).$ (17)
We simulate the dynamical equations on a $512\times 512$ descretized grid using finite differences to approximate the gradients. We use a vorticity stream function approach to solve for incompressibility. We non-dimensionalize our equations using the length scale $\epsilon=\sqrt{K_{Q}/\alpha_{Q}}=1$, and time scale $\tau=\gamma_{Q}/\alpha_{Q}=1$, setting our simulation size to $L=128$. We measure energy in units of the elastic constant, $K_{Q}=1$. The remaining parameters are given by $\rho=1$, $\eta=1$ and $\mu=16$. We enforce low Reynolds number by neglecting the advective term in Eq. 14. We simulate the system in the absence of flow alignment, $\lambda_{Q}=\lambda_{p}=0$. To ensure the polar field follows the nematic field closely, we set $K_{p}=K_{Q}/100$, $\alpha_{p}=\alpha_{Q}/25$ and $\alpha_{C}=4\alpha_{Q}/5$. Finally we fix the nematic active stress at $\zeta_{Q}=-0.2$ and the polar active force is taken from the range $\zeta_{P}\in[0.0,0.0026125]$.
We use this model to simulate specific creation and annihilation events, as this allows us to keep the grain boundary correlated with the orientation of the $^{1}/_{2}$ defect. The system is initialized with zero flow, $\underline{v}=0$. To simulate defect creation, we initialize the director fields with a localized bend distortion, see Extended Data Fig. 7a. If $\theta$ is the orientation of the polar field the bend distortion is given by
$\theta(\underline{r})=\frac{\pi}{2}+\exp\left(-\frac{r^{2}}{\sigma^{2}}\right)\sin\left(\frac{2\pi r_{y}}{L}\right)$ (18)
Here $\underline{r}=(r_{x},r_{y})$ is the position relative to the center of the simulation area of size $L$ and $\sigma=11L/128$ gives the scale of the distortion. For defect annihilation, we initialize the system with a pair of defects situated either side of the center of the simulation area, see Extended Data Fig. 7b. The field around a defect is given by
$\theta(\underline{r})=\frac{\phi_{1}-\phi_{2}}{2}.$ (19)
Here $\phi_{1}$ and $\phi_{2}$ are the polar angles around the locations of the $^{1}/_{2}$ and $^{1}/_{2}$ defects, respectively. The defects are located at $\underline{r}_{\mp}=(\pm\Delta,0)$ and $\Delta$ is half the separation between the defects; we set $\Delta=3L/32$. Finally, we reverse all polar vectors above the $x$ axis, which moves the grain boundary to the tail of the $^{1}/_{2}$ defect.
@@
As above we describe the nematic field by the tensor $\bm{Q}=S(\bm{n}\odot\bm{n}-\bm{I}/2)$ where $\bm{n}$ is a vector aligned with the nematic director. We now introduce the nematic orientation angle, $\psi$ which describes the average orientation of the nematic director at a point in space, hence
$\hat{\bm{n}}=[\cos(\psi),\sin(\psi)].$ (20)
Topological defects within the nematic field are characterized by a half integer winding number $k=\pm^{1}/_{2}$. The orientation of the nematic director field at a polar angle $\phi$ around a topological defect is given by:
@@
We describe the polar field with $\bm{p}$ and introduce the polar orientation angle, $\theta$, which is compatible with
$\hat{\bm{p}}=p[\cos(\theta),\sin(\theta)].$ (22)
For the following calculations we will neglect variations in the order parameter, i.e. $S=p=1$ and assume perfect correlation between the polar and nematic fields, i.e. $\bm{p}.\bm{n}=\pm 1$.
A half integer winding number, $k=\pm^{1}/_{2}$, is incompatible with the polar field $\bm{p}$, and necessitates the introduction of a discontinuous jump in the polar director angle, $\theta$. This is the grain boundary over which $\bm{{p}}$ goes through a $\pi$ rotation which begins and ends at a topological defect. We introduce the grain boundary as follows
@@
Flow field around a topological defect
At low Reynolds number, the flow caused by a force field $\bm{f}$ obeys the following equation
$\eta\nabla^{2}\bm{v}-\nabla P+\bm{f}=0.$ (25)
Where $\eta$ is the viscosity and $P$ is the pressure. For an incompressible fluid, solutions to this equation can be found by convoluting the force field $\bm{f}$ with the two dimensional Oseen tensor $\bm{G}$, hence
$v_{i}(\bm{r})=\int G_{ij}(\bm{r}-\bm{r}^{\prime})f_{j}(\bm{r}^{\prime})dA^{\prime},$ (26)
where the Oseen tensor is given by
$G_{ij}(\bm{r})=\frac{1}{4\pi\eta}\left[(\log{\frac{\mathcal{L}}{{r}}}-1)\delta_{ij}+\frac{r_{i}r_{j}}{r^{2}}\right]$ (27)
To calculate the flow around a defect, we assume that the active forces dominate, neglecting the passive stress tensor. In our model, the flow around a defect is dominated by the two active forces. The active force arising from the extensile active stress tensor is given by
$\bm{f}_{Q}=\zeta_{Q}\nabla.\bm{Q},$ (28)
and can be calculated as
$\bm{f}_{Q}=\frac{\zeta_{Q}}{2r}\bm{\hat{x}}$ (29)
where $r$ is the distance from the defect.
Following the calculation in Giomi et al. (2014) one can arrive at the following flow field arising from the active nematic forces
$\bm{v}^{Q}(r,\phi)=\frac{\zeta_{Q}}{12\eta}\left[3(R-r)+r\cos(2\phi),r\sin(2\phi)\right].$ (30)
Due to the linearity of Eq. 25, we can simply add this to the additional velocity caused by the polar body force.
The polar body force arising from self-propulsion is given by
$\bm{f}_{p}=\zeta_{p}\bm{p}.$ (31)
This can be calculated around a defect as
$\bm{f}_{p}=\zeta_{p}(-1)^{m}S(\phi,\phi_{g})\cos\left(\frac{\phi}{2}\right)\hat{\mathbf{x}}+\zeta_{p}(-1)^{m}S(\phi,\phi_{g})\sin\left(\frac{\phi}{2}\right)\hat{\mathbf{y}}$. (32)
Here we have introduced the sign function defined by
[…]
Following a similar process to that shown in Ref Giomi et al. (2014), one can arrive at a flow field that results from the active polar body force, which we denote $\bm{v}^{p}$. The equation representing $\bm{v}^{p}$ is very long and a full derivation is given in the appendix, along with the equivalent calculation for a $^{1}/_{2}$ defect. The final flow field around an active $^{1}/_{2}$ defect is given by $\bm{v}=\bm{v}^{Q}+\bm{v}^{p}$ and is shown in Fig. 5e.
To assess the active self-propulsion of a $^{1}/_{2}$ defect, we evaluate the generated velocity at the defect core, i.e. at $r=0$. This gives
$v(r=0)=\frac{R\zeta_{Q}}{4\pi\eta}\left[\pi+\frac{\zeta_{p}}{\zeta_{Q}}R\sin\left(\frac{\phi_{G}}{2}\right),-\frac{\zeta_{p}}{\zeta_{Q}}R\cos\left(\frac{\phi_{G}}{2}\right)\right].$ (34)
The argument of this velocity vector is displayed in Fig. 5f (evaluated for $R=1$ and $\eta=1$).
The torque on the defect is estimated from the curl of the velocity field at $r=0$, which is calculated numerically and displayed in Fig. 5g.
@@
Calculating the flow field around the defect is often a difficult process, however calculating the net force at the defect core is comparatively easy and gives very similar results. We include it here for completeness.
@@
$\underline{F}_{p}$ $=\oint\underline{f}_{p}(\phi)\textrm{d}\phi$ (35)
$\sim\zeta_{p}\oint\underline{\hat{p}}(\phi)\textrm{d}\phi$ (36)
@@
$\underline{F}$ $\sim\zeta_{p}\int_{-\pi}^{\phi_{g}}\underline{\hat{p}}(\phi)\textrm{d}\phi+\zeta_{p}\int_{\phi_{g}}^{\pi}\underline{\hat{p}}(\phi)\textrm{d}\phi$
$=\zeta_{p}\int_{-\pi}^{\phi_{g}}[\cos(k\phi+m\pi),\sin(k\phi+m\pi)]\textrm{d}\phi+$
$\zeta_{p}\int_{\phi_{g}}^{\pi}[\cos(k\phi+(m+1)\pi),\sin(k\phi+(m+1)\pi)]\textrm{d}\phi$
$=\zeta_{p}\left[\frac{[\sin(k\phi+m\pi),-\cos(k\phi+m\pi)]}{k}\right]_{-\pi}^{\phi_{g}}+$
$\zeta_{p}\left[\frac{[\sin(k\phi+(m+1)\pi),-\cos(k\phi+(m+1)\pi)]}{k}\right]_{\phi_{g}}^{\pi}$
$=\frac{2\zeta_{p}}{k}\left[\sin\left(k\phi_{g}+m\pi\right),-\cos\left(k\phi_{g}+m\pi\right)\right]$
@@
First, there is a net translational force on both $\pm^{1}/_{2}$ defects, meaning that we expect both defects to self-propel.
If $\phi_{g}=0$, as is the case immediately after creation, the additional force is perfectly perpendicular to the symmetry axis of the defect. When combined with the force from the extensile stress, this would imply a self-propulsion at an angle to the tail for $^{1}/_{2}$ defects.
@@
Finally, the force on the core of a $k=+^{1}/_{2}$ defect is perfectly aligned with the flow predicted by the Stokes flow calculation. If we add this to the net force arising from the extensile active stress, which is given by $\underline{F}_{Q}\sim[\zeta_{Q},0]$ we arrive at
$\underline{F}=\left[\zeta_{Q}+\zeta_{p}\sin\left(\phi_{g}/2\right),-\zeta_{p}\cos(\phi_{g}/2)\right],$ (38)
which shows the same behavior and symmetries as Eq. 34, see Extended Data Fig. 3c. Thus the self-propelled motion of the $^{1}/_{2}$ defects can be simply understood by considering the linear combination of the active forces at the core.
@@
We can assess the active torsional force around a defect caused by the polar active force by azimuthally integrating the azimuthal component of the polar force. This gives an equivalent picture to the predicted vorticity at the core of the defect. The curl of the polar force is given by:
@@
$\tau$ $\sim\zeta_{p}\oint_{-\pi}^{\pi}\sin(\theta-\phi)\textrm{d}\phi$ (41)
$=\zeta_{p}\oint_{-\pi}^{\pi}\sin((k-1)\phi+\pi*(H(\phi-\phi_{g})+m))\textrm{d}\phi$ (42)
$=\zeta_{p}\left[\frac{-\cos((k-1)\phi+m\pi)}{k-1}\right]_{-\pi}^{\phi_{g}}+$ (43)
$\zeta_{p}\left[\frac{-\cos((k-1)\phi+(m+1)\pi)}{k-1}\right]_{\phi_{g}}^{\pi}$ (44)
$=\frac{\zeta_{p}}{k-1}\left[2\cos((k-1)\phi_{g}+(m+1)\pi)+\right$. (45)
@@
Second, the magnitude of the torque depends on the orientation of the grain boundary. The torque disappears when $(k-1)\phi_{g}+m\pi=(n+0.5)\pi$, $n\in\mathbb{Z}$. This is equivalent to when the net additional net force aligns with an axis of mirror symmetry.
@@
Finally, since the curl of the force arising from the active stress is zero around the core of a defect, any net vorticity around the defect must be arise from the curl of the body force. We plot $\tau$ around a $^{1}/_{2}$ defect as a function of $\phi_{g}$ and $\zeta_{p}$ in Extended Data Fig. 3d, and we see that it is a good predictor for the vorticity around a defect (Fig. 5g.